# Single-file dynamics with general charge measures

### v2

@@ Abstract
keywords:
charged single-file dynamics, anomalous fluctuations, kinetic constraints, dynamical phase transitions
1 Introduction
@@
One of the most widely applicable results of probability theory is the central limit theorem which, in its most basic form, characterizes fluctuations of the average of $n$ independent identically distributed variables sampled from a distribution with finite variance. The theorem states that typical fluctuations around the mean value are of order $\mathcal{O}(n^{1/2})$ and their distribution is asymptotically Gaussian. The central limit theorem has been further refined to cover the important cases of not identically distributed or weakly correlated variables [1].
@@
A natural extension to the study of typical fluctuations is given by large deviation theory, characterizing large fluctuations, of order $\mathcal{O}(n)$, away from the mean. A basic problem of large deviation theory is to establish a large deviation principle for tail probabilities, i.e. to show that large fluctuations are suppressed exponentially in $n$ and to characterize the rate of suppression, encapsulated in the so-called rate function. The central limit theorem can then often be understood as the quadratic expansion of the rate function around the mean value.
@@
The theory of large deviations is the natural language to describe the central notion of concentration of measure in statistical mechanics both in and out of equilibrium [2, 3]. Fluctuations of conserved quantities in many-body systems are captured by the so-called full-counting statistic of corresponding time-integrated current densities, a dynamical analogue of the partition function in equilibrium statistical mechanics. A time-integrated current is the many-body analogue of the sum of a sequence of variables, with the proviso that the current densities at different times are now correlated by the dynamics with randomness necessarily entering even for deterministic dynamics due to sampling from an ensemble of initial conditions.
@@
This naturally leads to the problem of characterizing current fluctuations on both the large and typical scale. The question of large fluctuations in ergodic systems with a single conserved quantity has been thoroughly addressed by the development of macroscopic fluctuations theory [4]. The theory establishes a large deviation principle and formulates a variational problem, the solution of which recovers the rate function. Recently, building upon the theory of generalized hydrodynamics [5, 6], an analogous ballistic macroscopic fluctuations theory [7] has been developed to describe large fluctuations in integrable systems. Similar results have also been derived by mapping the full-counting statistics in a non-equilibrium bipartite ensemble to a dual equilibrium problem using space-time duality [8].
@@
Establishing a large deviation principle quantifies the structure of large fluctuations and its expansion to quadratic order suggest asymptotic Gaussianity of typical fluctuations. This might lead to the erroneous conclusion that a large deviation principle implies Gaussianity on the typical scale. However, inferring behavior on the typical scale from the large scale involves a rescaling for which the logarithm of the full-counting statistics needs to satisfy an additional regularity condition [9]. While Gaussianity of typical fluctuations in a class of chaotic quantum systems has been recently demonstrated [10], numerical studies of classical integrable spin chains [11, 12] have found clear violations of asymptotic Gaussianity in equilibrium ensembles at half-filling.
@@
This unexpected observation has led to the study and exact solution of equilibrium fluctuations [13] in a simple interacting classical cellular automaton of impenetrable ballistic $\mathbb{Z}_{2}$-charged particles first studied in [14]. Typical fluctuations were found to follow an M-Wright distribution, normally associated with fractional diffusion [15, 16, 17, 18]. The violation of Gaussianity was traced to a complex zero of the full-counting statistics colliding with the origin, breaking the regularity condition in a process reminiscent of an equilibrium phase transition. Shortly thereafter, the same distribution was derived in the low temperature regime of the sine-Gordon model [19], see also [20] for an earlier derivation in the same setting using form factors. It was soon realized that the observed phenomenology is closely related to fragmentation of the configuration space by a strict kinetic constraint, giving rise to the class of charged single-file dynamics whose fluctuations were studied in detail in [21]. Besides anomalous fluctuations in equilibrium, the non-equilibrium setting gave rise to a rich interplay of dynamical phase transitions in the rate function of large fluctuations. Apart from fluctuations, single-file kinetic constraints have also been used to study dynamicala correlation functions at hydrodynamics scales [22] and in some exact solutions [23, 24, 25].
@@
In Section 2, we introduce the notion of single-file dynamics and define the family of initial bipartite ensembles studied in the work. Attaching charge degrees of freedom to particles, we introduce inert charged single-file dynamics and comment on the role of charge inertness. We end the introductory section by giving a brief overview of the basic tools that we use in the remainder of the work.
In Section 3 we derive the exact finite-time distribution of charge-particle fluctuations by dressing particle fluctuations by means of a conditional probability distribution most easily expressed in Fourier space. The finite-time dressing is shown to be equivalent to a simple substitution rule acting on the particle full-counting statistics.
In Section 4 we study the asymptotic behavior of the dressing. Starting with typical fluctuations, we show that they generically depend only on the first two moments of the charge measure. With mild technical assumption on particle fluctuations, we show that typical fluctuations in generic bipartite initial ensembles are asymptotically Gaussian with a variance that decomposes into a contribution from the charge mean and another from the variance of the charge measure. In equilibrium ensembles with vanishing mean charge we instead find that typical fluctuations are parametrically suppressed and converge to an M-Wright distribution.
Turning our attention to large fluctuations, we derive an integral representation of the conditional charge-particle rate function and a corresponding saddle-point equation. For charge measures with rapidly decaying tails we show that the saddle-point equation has exactly one solution on the imaginary axis. We formulate a two-step minimization problem for the charge rate function. Properties of the saddle-point equations allow us to establish convexity of the optimization problem which in turn enables a general description of the charge rate function. In particular, we show that the charge rate function can exhibit dynamical phase transitions of first and second order.
We conclude in Section 5 by summarizing our results and putting them in the context of recent advances in the study of fluctuations in integrable systems and related models.
2 Charged single-file dynamics
2.1 Particle dynamics
@@
Consider a system of $Q_{p}$ classical particles on a periodic ring of length $L$ and label particle positions in configuration space $\mathcal{Q}_{p}\equiv[-L/2,L/2)^{Q_{p}}$ at time $t$ as ${{\mathbf{x}}}(t)\equiv(x_{1}(t),\ldots,x_{Q_{p}}(t))\in\mathcal{Q}_{p}$, where we identify $x_{j}=x_{j}+L$. By convention, assuming a generic position (one in which no two particles overlap $x_{i}(t_{0})\neq x_{j}(t_{0}),\ \forall i,j\in\mathbb{N}_{Q_{p}}$), we order the particles from left to right at initial time $t_{0}$
$x_{1}(t_{0})<x_{2}(t_{0})<\ldots<x_{Q_{p}}(t_{0}).$ (1)
The particles’ dynamics is given by a map $\phi_{s}:\mathcal{Q}_{p}\to\mathcal{Q}_{p}$ which propagates the particles’ positions forward in time
$\phi_{s}\left[{\mathbf{x}}(t)\right]={\mathbf{x}}(t+s).$ (2)
@@
We consider dynamics $\phi_{s}$ that admit a flat invariant measure characterized by the average particle density $\rho$. We specialize to a simple nonequilibrium setting by considering a bipartite initial measure consisting of two equilibrium measures joined at the origin with particle densities $\rho_{\pm}$ to the right/left of the origin respectively
$\varrho_{L}({\mathbf{x}},t_{0})=\prod_{j=1}^{Q_{p}}p_{L}(x_{j}),\qquad p_{L}(x)=\frac{2L^{-1}}{\rho_{-}+\rho_{+}}\begin{cases}\rho_{-},&-L/2\leq x<0,\\ \rho_{+},&0\leq x<L/2.\end{cases}$ (3)
@@
The integral of the local particle density $\rho_{p}(x,t)\equiv\sum_{n=1}^{Q_{p}}\delta(x-x_{n}(t))$ gives the total particle number $Q_{p}=\int_{-L/2}^{L/2}{\rm d}x\,\rho_{p}(x,t)$. In the remainder of the paper we consider systems in the thermodynamics limit at finite particle density with the corresponding initial measure
$\varrho({\mathbf{x}},t_{0})=\lim_{\stackrel{{\scriptstyle Q_{p},L\to\infty}}{{Q_{p}/L=(\rho_{-}+\rho_{+})/2}}}\varrho_{L}({\mathbf{x}},t_{0}).$ (4)
2.2 Inert charged dynamics
@@
We attach a scalar charge $c\in\mathcal{C}\subseteq\mathbb{R}$ to each particle, sampled from a normalized measure, $\omega_{\mu}:\mathcal{C}\to\mathbb{R}^{+}_{0}$, $\int_{\mathcal{C}}{\rm d}\omega_{\mu}=1$, with a finite mean $|\int_{\mathcal{C}}{\rm d}\omega_{\mu}\,c|=|\mu|<\infty$ and variance $\int_{\mathcal{C}}{\rm d}\omega_{\mu}(c-\mu)^{2}=\sigma^{2}_{\omega}<\infty$. We consider measures whose support is not sign-definite, ${\rm s}_{\pm}\equiv{\rm supp}(\omega_{\mu})\cap\mathbb{R^{\pm}}\neq\emptyset$ and denote ±$\pm c_{\pm}\equiv\displaystyle\max_{c\in{\rm s}_{\pm}}\pm c$.
@@
The enlarged configuration space $\mathcal{Q}^{c}_{p}\equiv\mathcal{Q}_{p}\times\mathcal{Q}_{c}$ is spanned by $({\mathbf{x}}(t),{\mathbf{c}})$, where ${\mathbf{c}}\equiv(c_{1},\ldots,c_{Q_{p}})\in\mathcal{Q}_{c}\equiv\mathcal{C}^{Q_{p}}$ and $c_{j}$ is the charge of the $j$-th particle. We consider inert charges, meaning that the dynamics of the charged system $\phi_{s}^{c}:\mathcal{Q}^{c}_{p}\to\mathcal{Q}^{c}_{p}$ decomposes as
$\phi_{s}^{c}[\left({\mathbf{x}}(t),{\mathbf{c}}\right)]=\left(\phi_{s}[{\mathbf{x}}(t)],{\mathbf{c}}\right)=\left({\mathbf{x}}(t+s),{\mathbf{c}}\right).$ (5)
@@
$\varrho^{c}_{L}({\mathbf{x}},{\mathbf{c}},t_{0})=\prod_{j=1}^{Q_{p}}p_{L}(x_{j})\omega_{\mu_{s_{j}}}(c_{j}),$ (6)
where $s_{j}=\textrm{sgn}(x_{j})$ and denote its thermodynamic limit as
$\varrho^{c}({\mathbf{x}},{\mathbf{c}},t_{0})=\lim_{\stackrel{{\scriptstyle Q_{p},L\to\infty}}{{Q_{p}/L=(\rho_{-}+\rho_{+})/2}}}\varrho^{c}_{L}({\mathbf{x}},{\mathbf{c}},t_{0}).$ (7)
2.3 Inert charged single-file dynamics
We obtain inert charged single-file dynamics by imposing a strict kinetic constraint on the composite dynamics $\phi_{s}^{c}$. In particular, we consider single-file dynamics for which the initial particle ordering (1) is preserved by the dynamics $\phi_{s}^{c}$
@@ 2.3 Inert charged single-file dynamics
which, due to charge inertness (5), is the same as imposing the single-file constraint on the particle dynamics $\phi_{s}$. We note that in an arbitrary crossing (i.e. one violating the single-file condition (8)) particle dynamics it is always possible to enforce the single-file constraint by adding an infinitely strong repulsive hardcore point interaction to the particles. For momentum conserving particle dynamics this leads to elastic particle scattering and reduces to permuting particle labels of colliding particles.
2.4 Full-counting statistics
Conservation of particle number $Q_{p}$ and total charge $Q_{c}$ means that their local densities satisfy continuity equations
$\partial_{t}\rho_{i}(x,t)+\partial_{x}j_{i}(x,t)=0,$ (9)
where $i\in\{p,c\}$ and $j_{i}(x,t)$ is the local current density. The corresponding integrated currents at the origin
@@ 2.4 Full-counting statistics
where $\lambda_{p}$ and $\lambda_{c}$ are the particle and charge counting field respectively and the average $\langle\bullet\rangle_{\mathcal{Q}_{p}^{c}}$ is taken with respect to the initial nonequilibrium measure (7). The corresponding joint probability distribution $\mathcal{P}_{c,p}(J_{c},J_{p}|t)$ is recovered by the inverse Laplace transform of the full-counting statistics
@@
where $\mathcal{L}_{x}[f(x)](y)\equiv\int_{-\infty}^{\infty}{\rm d}x\,e^{-yx}f(x)$ is the bilateral Laplace transform with respect to $x$. The univariate probability distributions $\mathcal{P}_{i}(J_{i}|t)$ of the integrated particle/charge current are obtained by marginalizing the joint probability distribution. At the level of full-counting statistics marginalization amounts to setting a counting field to zero
$G_{i}(\lambda_{i}|t)=G_{c,p}(\lambda_{c},\lambda_{p}|t)|_{\lambda_{j\neq i}=0},\qquad i,j\in\{c,p\}.$ (13)
The charge full-counting statistics (13) involves averaging over the composite initial measure (7) as charges are bound to particles. However, since inert charge do not effect particle dynamics, the inert particle full-counting statistics can be restricted to averaging solely over the particle initial measure (4)
@@
Noting that the integrated particle current $J_{p}$ is invariant under arbitrary permutations of particle labels, the particle full-counting statistics of a single-file dynamics with elastic scattering is identical to that of a crossing dynamics by considering quasi-particles instead of particles as e.g. in the hard-rod gas or generalized hydrodynamics [26]. In particular, even a free particle dynamics, whose particle full-counting statistics is trivially related to the one-particle full-counting statistics, gives an interacting dynamics upon imposition of the single-file condition, which was used in [21, 13] to derive the exact finite-time particle and charge full-counting statistics for $\mathbb{Z}_{2}$ charges.
3 Dressing
We now consider the following question:
Given a probability distribution of the integrated particle current $\mathcal{P}_{p}(J_{p}|t)$ of a single-file particle dynamics $\phi_{s}$, what is the probability distribution of the integrated charge current $\mathcal{P}_{c}(J_{c}|t)$ of the inert charged single-file dynamics $\phi_{s}^{c}$?
Fluctuations of particles and charges are easily related in terms of their probability distributions. The joint charge-particle distribution $\mathcal{P}_{c,p}$ factorizes into a product of the particle distribution and a time-independent conditional probability $\mathcal{P}_{c|p}$
$\mathcal{P}_{c,p}({J}_{c},J_{p}|t)=\mathcal{P}_{c|p}({J}_{c}|{J}_{p})\mathcal{P}_{p}({J}_{p}|t).$ (15)
The conditional distribution gives the probability of the integrated charge current $J_{c}$ conditioned on a value of the integrated particle current $J_{p}$. Given that inert charges do not influence particle dynamics, the conditional probability takes the form

### v3

@@ Abstract
I Introduction
@@
One of the most widely applicable results of probability theory is the central limit theorem that, in its basic form, characterizes fluctuations of the average of $n$ independent identically distributed variables sampled from a distribution with finite variance. The theorem states that typical fluctuations around the mean value are of order $\mathcal{O}(n^{1/2})$ and their distribution is asymptotically Gaussian. The central limit theorem has been further refined to cover the important cases of not identically distributed or weakly correlated variables [1].
@@
A natural extension to the study of typical fluctuations is given by large deviation theory, characterizing large fluctuations of order $\mathcal{O}(n)$ away from the mean. A basic problem of large deviation theory is to establish a large deviation principle for tail probabilities, i.e. to show that large fluctuations are suppressed exponentially in $n$ and to characterize the rate of suppression, encapsulated in the rate function. The central limit theorem can then often be understood as the quadratic expansion of the rate function around the mean value.
@@
The theory of large deviations is also the natural language to describe the central notion of concentration of measure in statistical mechanics both in and out of equilibrium [2, 3]. Fluctuations of conserved quantities in many-body systems are captured by the full-counting statistic of corresponding time-integrated current densities, the time-integrated current being the many-body analogue of the sum of a sequence of variables, with the proviso that current densities at different times are now correlated by the dynamics with randomness necessarily entering even for deterministic dynamics due to sampling from an ensemble of initial conditions.
@@
This naturally leads to the problem of characterizing current fluctuations on both the large and typical scale. The question of large fluctuations in ergodic systems with a single conserved quantity has been thoroughly addressed by the development of macroscopic fluctuations theory, see [4] for a review. The theory establishes a large deviation principle for fluctuations of the integrated current and formulates a variational problem, the solution of which recovers the rate function. Recently, building upon the theory of generalized hydrodynamics [5, 6], an analogous ballistic macroscopic fluctuations theory [7] has been developed to describe large fluctuations in integrable systems. Similar results have also been derived by mapping the full-counting statistics in a non-equilibrium bipartite ensemble to a dual equilibrium problem using space-time duality [8].
@@
Establishing a large deviation principle quantifies the structure of large fluctuations and its expansion to quadratic order suggest asymptotic Gaussianity of typical fluctuations. This might lead to the erroneous conclusion that a large deviation principle directly implies Gaussianity on the typical scale. However, inferring behavior on the typical scale from the large scale involves a rescaling for which the logarithm of the full-counting statistics needs to satisfy an additional regularity condition [9]. While Gaussianity of typical fluctuations in a class of chaotic quantum systems has been recently demonstrated [10], numerical studies of classical integrable spin chains [11, 12] have found clear violations of asymptotic Gaussianity in equilibrium ensembles at half-filling.
@@
This unexpected observation has led to an exact solution of equilibrium fluctuations [13] in a simple interacting classical cellular automaton of impenetrable ballistic $\mathbb{Z}_{2}$-charged particles first studied in [14]. Typical fluctuations were found to follow an M-Wright distribution, normally associated with fractional diffusion [15, 16, 17, 18]. The violation of Gaussianity was traced to a complex zero of the full-counting statistics colliding with the origin, breaking the regularity condition in a process reminiscent of an equilibrium phase transition. Shortly thereafter, the same distribution was derived in the low temperature regime of the sine-Gordon model [19], see also [20] for an earlier derivation in the same setting using form factors. It was soon realized that the observed phenomenology is closely related to fragmentation of the configuration space by a strict kinetic constraint, giving rise to the class of charged single-file dynamics whose fluctuations were studied in detail in [21]. Besides anomalous fluctuations in equilibrium, the non-equilibrium setting gave rise to a rich interplay of dynamical phase transitions in the rate function of large fluctuations. Apart from fluctuations, single-file kinetic constraints have also been used to study dynamical correlation functions at hydrodynamics scales [22] and in some exact solutions [23, 24, 25].
@@
The paper is organized as follows: In Section II, we introduce single-file dynamics and the basic concepts used in the remainder of the work. In Section III we derive the exact finite-time joint charge-particle fluctuations in terms of particle fluctuations. In Section IV we study the time-asymptotic behavior of joint fluctuations at both the typical and large scale. We conclude in Section V by putting our results in the context of recent advances in the study of fluctuations in integrable systems and related models.
II Charged single-file dynamics
@@
Consider a system of $Q_{p}$ classical particles on a periodic ring of length $L$ and label particle positions in configuration space $\mathcal{Q}_{p}\equiv[-L/2,L/2)^{Q_{p}}$ at time $t$ as ${\bf x}(t)\equiv(x_{1}(t),\ldots,x_{Q_{p}}(t))\in\mathcal{Q}_{p}$, where we identify $x_{j}=x_{j}+L$. Assuming a generic position (one in which no two particles coincide), we order the particles from left to right at initial time $t_{0}$, $x_{1}(t_{0})<x_{2}(t_{0})<\ldots<x_{Q_{p}}(t_{0}).$ The particles’ dynamics is given by a map $\phi_{s}:\mathcal{Q}_{p}\to\mathcal{Q}_{p}$ which propagates the particles’ positions forward in time
$\phi_{s}\left[{\bf x}(t)\right]={\bf x}(t+s).$ (1)
@@
We consider dynamics $\phi_{s}$ that admit a flat invariant measure characterized by the average particle density $\rho$ and specialize to a simple nonequilibrium setting by considering a bipartite initial measure consisting of two equilibrium measures joined at the origin with particle densities $\rho_{\pm}$ to the right/left of the origin respectively
$\varrho_{L}({\bf x},t_{0})=\prod_{j=1}^{Q_{p}}p_{L}(x_{j}),$ (2)
with the one particle density
$p_{L}(x)=\frac{2L^{-1}}{\rho_{-}+\rho_{+}}\begin{cases}\rho_{-},&-L/2\leq x<0,\\ \rho_{+},&0\leq x<L/2.\end{cases}$ (3)
@@
The integral of the local particle density $\rho_{p}(x,t)\equiv\sum_{n=1}^{Q_{p}}\delta(x-x_{n}(t))$ gives the total particle number $Q_{p}=\int_{-L/2}^{L/2}{\rm d}x\,\rho_{p}(x,t)$. We consider systems in the thermodynamics limit at finite particle density with the corresponding initial measure
$\varrho({\bf x},t_{0})=\lim_{\stackrel{{\scriptstyle Q_{p},L\to\infty}}{{Q_{p}/L=(\rho_{-}+\rho_{+})/2}}}\varrho_{L}({\bf x},t_{0}).$ (4)
Figure 1: Many-body trajectory of an inert charged single-file system consisting of ballistically propagating hardcore particles (broken straight lines) carrying inert charges (blue to red) sampled from a Gaussian charge measure (17).
@@
We attach a scalar charge $c\in\mathcal{C}\subseteq\mathbb{R}$ to each particle, sampled from a normalized measure, $\omega_{\mu}:\mathcal{C}\to\mathbb{R}^{+}_{0}$, $\int_{\mathcal{C}}{\rm d}\omega_{\mu}=1$, with a finite mean $|\int_{\mathcal{C}}{\rm d}\omega_{\mu}\,c|=|\mu|<\infty$ and variance $\int_{\mathcal{C}}{\rm d}\omega_{\mu}(c-\mu)^{2}=\sigma^{2}_{\omega}<\infty$ whose support is not sign-definite, ${\rm s}_{\pm}\equiv{\rm supp}(\omega_{\mu})\cap\mathbb{R^{\pm}}\neq\emptyset$ and denote ±$\pm c_{\pm}\equiv\displaystyle\max_{c\in{\rm s}_{\pm}}\pm c$.
@@
The enlarged configuration space $\mathcal{Q}^{c}_{p}\equiv\mathcal{Q}_{p}\times\mathcal{Q}_{c}$ is spanned by $({\bf x}(t),{\bf c})$, where ${\bf c}\equiv(c_{1},\ldots,c_{Q_{p}})\in\mathcal{Q}_{c}\equiv\mathcal{C}^{Q_{p}}$ and $c_{j}$ is the charge of the $j$-th particle. We consider inert charges, meaning that the dynamics of the charged system $\phi_{s}^{c}:\mathcal{Q}^{c}_{p}\to\mathcal{Q}^{c}_{p}$ decomposes as
$\phi_{s}^{c}[\left({\bf x}(t),{\bf c}\right)]=\left(\phi_{s}[{\bf x}(t)],{\bf c}\right)=\left({\bf x}(t+s),{\bf c}\right).$ (5)
@@
$\varrho^{c}_{L}({\bf x},{\bf c},t_{0})=\prod_{j=1}^{Q_{p}}p_{L}(x_{j})\omega_{\mu_{s_{j}}}(c_{j}),$ (6)
where $s_{j}=\textrm{sgn}(x_{j})$ and take the thermodynamic limit
$\varrho^{c}({\bf x},{\bf c},t_{0})=\lim_{\stackrel{{\scriptstyle Q_{p},L\to\infty}}{{Q_{p}/L=(\rho_{-}+\rho_{+})/2}}}\varrho^{c}_{L}({\bf x},{\bf c},t_{0}).$ (7)
We further consider single-file dynamics for which the initial particle ordering is preserved by the dynamics $\phi_{s}^{c}$
@@
which, due to charge inertness (5), is the same as imposing the single-file constraint on the particle dynamics $\phi_{s}$. A snapshot of an inert charged single-file system in the thermodynamic limit is shown in Figure 1. Note that in an arbitrary crossing (i.e. one violating the single-file condition (8)) particle dynamics it is always possible to enforce the single-file constraint by adding a strong repulsive hardcore interaction to the particles.
II.1 Full-counting statistics
Conservation of particle number $Q_{p}$ and total charge $Q_{c}$ means that their local densities satisfy continuity equations $\partial_{t}\rho_{i}(x,t)+\partial_{x}j_{i}(x,t)=0$, where $i\in\{p,c\}$ and $j_{i}(x,t)$ is the local current density. The corresponding integrated currents at the origin
@@ II.1 Full-counting statistics
where $\lambda_{p}$ and $\lambda_{c}$ are the particle and charge counting fields respectively and the average $\langle\bullet\rangle_{\mathcal{Q}_{p}^{c}}$ is taken with respect to the initial nonequilibrium measure (7). The corresponding joint probability distribution $\mathcal{P}_{c,p}(J_{c},J_{p}|t)$ is recovered by the inverse Laplace transform of the full-counting statistics
@@
where $\mathcal{L}_{x}[f(x)](y)\equiv\int_{-\infty}^{\infty}{\rm d}x\,e^{-yx}f(x)$ is the bilateral Laplace transform with respect to $x$. The univariate probability distributions $\mathcal{P}_{i}(J_{i}|t)$ of the integrated particle/charge current are obtained by marginalizing the joint probability distribution, which at the level of full-counting statistics amounts to setting a counting field to zero, $G_{i}(\lambda_{i}|t)=G_{c,p}(\lambda_{c},\lambda_{p}|t)|_{\lambda_{j\neq i}=0}$. The charge full-counting statistics $G_{c}$ involves averaging over the composite initial measure (7) as charges are bound to particles. However, since inert charge do not effect particle dynamics, the particle full-counting statistics involves averaging solely over the particle initial measure (4)
@@
Noting that the integrated particle current $J_{p}$ is invariant under arbitrary permutations of particle labels, the particle full-counting statistics of a single-file dynamics with elastic scattering is identical to that of a crossing dynamics by considering quasi-particles (unbroken straight lines in Fig. 1) instead of particles as e.g. in the hard-rod gas or generalized hydrodynamics [26]. In particular, even a free particle dynamics gives an interacting dynamics upon imposition of the single-file condition, which was used in [21, 13] to derive the exact finite-time particle and charge full-counting statistics for $\mathbb{Z}_{2}$ charges.
III Dressing
Charge current fluctuations can be obtained from fluctuations of the particle current by a combinatorial dressing procedure due to charge inertness (5) as we discuss below. The general form of the dressing simplifies for single-file dynamics (8), resulting in a tractable integral representation.
The joint charge-particle distribution $\mathcal{P}_{c,p}$ of an arbitrary (i.e. not necessarily inert or single-file) charged particle dynamics can be factorized into a product of the particle distribution and a time-independent conditional probability, $\mathcal{P}_{c,p}({J}_{c},J_{p}|t)=\mathcal{P}_{c|p}({J}_{c}|{J}_{p})\mathcal{P}_{p}({J}_{p}|t).$ The conditional distribution gives the probability of the integrated charge current $J_{c}$ conditioned on a value of the integrated particle current $J_{p}$. For a particle dynamics with inert charges the conditional probability is given by